# Finite-Sample Analysis of Policy Evaluation for Robust Average Reward Reinforcement Learning

### v2

@@ 2.2 Robust Bellman Operator
Motivated by Theorem 1, we define the robust Bellman operator, which forms the basis for our policy evaluation procedure.
Definition 1 (Robust Bellman Operator, Wang et al. (2023b)).
@@
The operator $\mathbf{T}_{g}$ transforms a value function $V$ by incorporating the worst-case transition effect. A key challenge in solving the robust Bellman equation is that $\mathbf{T}_{g}$ does not satisfy contraction under standard norms, preventing the use of conventional fixed-point iteration. To cope with this problem, we establish that $\mathbf{T}_{g}$ is a contraction under some constructed semi-norm. This allows us to develop provably efficient stochastic approximation algorithms. Throughout this paper, we make the following standard assumption regarding the structure of the induced Markov chain.
Assumption 1.
The Markov chain induced by $\pi$ is irreducible and aperiodic for all $\mathsf{P}\in\mathcal{P}$.
Assumption 1 is used widely in the current robust average reinforcement learning literature Wang et al. (2023a, b, 2024); Sun et al. (2024). This assumption ensures that, under any transition model within the uncertainty set, the policy $\pi$ induces a single recurrent communicating class. A well-known result in average-reward MDPs states that under Assumption 1, the average reward is independent of the starting state, i.e., for any $\mathsf{P}\in\mathcal{P}$ and all $s,s^{\prime}\in\mathcal{S}$, we have $g^{\pi}_{\mathsf{P}}(s)=g^{\pi}_{\mathsf{P}}(s^{\prime})$. Thus, we can drop the dependence on the initial state and simply write $g^{\pi}_{\mathsf{P}}$ as the robust average reward.
Remark 1.
We note that when the radius of uncertainty sets are small enough, Assumption 1 could be relaxed into assuming only the nominal transition being irreducible and aperiodic. We provide the detailed discussions in Appendix B.
@@ 3 Semi-Norm Contraction of Robust Bellman Operators
Under Assumption 1, we are able to establish the semi-norm contraction property. For motivation, we first establish the semi-norm contraction property of the non-robust average reward Bellman operator for a policy $\pi$ under transition $\mathsf{P}$ defined as follows:
@@
Under ergodicity, the one‐step transition matrix (denoted as $\mathsf{P}^{\pi}$) has a unique stationary distribution $d^{\pi}$, define the stationary projector $E=\mathbf{e}^{\top}d^{\pi}$, then the fluctuation matrix (defined as $Q^{\pi}=\mathsf{P}^{\pi}-E$) has all eigenvalues strictly inside the unit circle. Standard finite‐dimensional theory (via the discrete Lyapunov equation Horn and Johnson (2012)) would produce a norm $\|\cdot\|_{Q}$ on $\mathbb{R}^{S}$ such that there is a constant $\alpha\in(0,1)$ such that for any $x\in\mathbb{R}^{S}$, $\|Q^{\pi}x\|_{Q}\leq\alpha\|x\|_{Q}$. We then build the semi-norm as follows:
@@
The concrete proof of Lemma 1 including the detailed construction of the semi-norm $\|\cdot\|_{\mathsf{P}}$ is in Appendix C.1, where the properties of irreducible and aperiodic finite state Markov chain are utilized. Thus, we show the (non-robust) average reward Bellman operator $\mathbf{T}_{g}^{\mathsf{P}}$ is a strict contraction under $\|\cdot\|_{\mathsf{P}}$. Based on the above motivations, we now formally establish the contraction property of the robust average reward Bellman operator by leveraging Lemma 1 and the compactness of the uncertainty sets.
Theorem 2.
Under Assumption 1, and if $\mathcal{P}$ is compact, there exists a semi-norm $\|\cdot\|_{\mathcal{P}}$ with kernel $\{c\mathbf{e}:c\in\mathbb{R}\}$ such that the robust Bellman operator $\mathbf{T}_{g}$ is a contraction. Specifically, there exist $\gamma\in(0,1)$ such that
@@
For any $\mathsf{P}\in\mathcal{P}$, the one‐step transition matrix $\mathsf{P}^{\pi}$ has a unique stationary projector $E_{\mathsf{P}}$ due to ergodicity. Since $\mathcal{P}$ is compact, the family of fluctuation matrices $\{Q_{\mathsf{P}}^{\pi}=\mathsf{P}^{\pi}-E_{\mathsf{P}}:\mathsf{P}\in\mathcal{P}\}$ has joint spectral radius strictly less than $1$. By Lemma 10in Berger and Wang (1992), one is able to construct an “extremal norm” (denoted as $\|\cdot\|_{\rm ext}$) under which every $Q_{\mathsf{P}}^{\pi}$ contracts by a uniform factor $\alpha\in(0,1)$. Mimicking the non-robust case in Lemma 1, we similarly define
@@
The concrete proof of Theorem 2 along with the detailed construction of the semi-norm $\|\cdot\|_{\mathcal{P}}$ are in Appendix C.2. Since all the uncertainty sets listed in Section 2.1 are closed and bounded in a real vector space, these uncertainty sets are all compact and satisfy the contraction property in Theorem 2.
@@ 4 Efficient Estimators for Uncertainty Sets
where the learning rate $\eta_{t}$ would be specified in Section 5. The detailed analysis and complexities of the general stochastic approximation in the form of (17) is provided in Appendix D. Theorem 8 implies that if $\hat{\mathbf{T}}_{g}(V)$, being an estimator of $\mathbf{T}_{g}(V)$, could be constructed with bounded variance and small bias, $V_{t}$ converges to a solution of the Bellman equation in (4). However, the challenge of constructing our desired $\hat{\mathbf{T}}_{g}(V)$ lies in the construction of the support function estimator $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$.
In this section, we aim to construct an estimator $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ for all $s\in\mathcal{S}$ and $a\in\mathcal{A}$ in various uncertainty sets. Recall that the support function ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$ represents the worst-case transition effect over the uncertainty set $\mathcal{P}^{a}_{s}$ as defined in the robust Bellman equation in Theorem 1. The explicit forms of ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$ for different uncertainty sets were characterized in (6)-(9). Our goal in this section is to construct efficient estimators $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ that approximates ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$ while maintaining controlled variance and finite sample complexity.
@@
Recall that the $\delta$-contamination uncertainty set is $\mathcal{P}^{a}_{s}=\{(1-\delta)\tilde{\mathsf{P}}^{a}_{s}+\delta q:q\in\Delta(\mathcal{S})\},$ where $10<\delta<$ 1 is the radius. Since the support function can be computed by (6) and the expression is linear in the nominal transition kernel $\tilde{\mathsf{P}}^{a}_{s}$. A direct approach is to use the transition to the subsequent state to construct our estimator:
@@
where $s^{\prime}$ is a subsequent state sample after $(s,a)$. Hence, the sample complexity of (18) is just one. Lemma 12 from Wang et al. (2023b) states that $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ obtained by (18) is unbiased and has bounded variance as follows:
@@
Regarding TV and Wasserstein distance uncertainty sets, they have a nonlinear relationship between the nominal distribution $\tilde{\mathsf{P}}^{a}_{s}$ and the support function ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$. Previous works such as (Blanchet and Glynn, 2015; Blanchet et al., 2019; Wang et al., 2023b) have proposed a multi-level Monte-Carlo (MLMC) method for obtaining an unbiased estimator of ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$ with bounded variance. However, their approaches require drawing $2^{N+1}$ samples where $N$ is sampled from a geometric distribution $\mathrm{Geom}(\Psi)$ with parameter $\Psi\in(0,0.5)$. This operation would need infinite samples in expectation for obtaining each single estimator as $\mathbb{E}[2^{N+1}]=\sum^{\infty}_{N=0}2^{N+1}\Psi(1-\Psi)^{N}=\sum^{\infty}_{N=0}2\Psi(2-2\Psi)^{N}\rightarrow\infty$. To handle the above problem, we aim to provide an estimator $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ with finite sample complexity and small enough bias. We construct a truncated-MLMC estimator under geometric sampling with parameter $\Psi=0.5$ as shown in Algorithm 1.
@@
4:$\hat{\mathsf{P}}^{a,E}_{s,N^{\prime}+1}\leftarrow\frac{1}{2^{N^{\prime}}}\sum_{i=1}^{2^{N^{\prime}}}\mathbbm{1}_{\{s^{\prime}_{2i}\}}$
5:$\hat{\mathsf{P}}^{a,O}_{s,N^{\prime}+1}\leftarrow\frac{1}{2^{N^{\prime}}}\sum_{i=1}^{2^{N^{\prime}}}\mathbbm{1}_{\{s^{\prime}_{2i-1}\}}$
6:$\hat{\mathsf{P}}^{a}_{s,N^{\prime}+1}\leftarrow\frac{1}{2^{N^{\prime}+1}}\sum_{i=1}^{2^{N^{\prime}+1}}\mathbbm{1}_{\{s^{\prime}_{i}\}}$
7:$\hat{\mathsf{P}}^{a,1}_{s,N^{\prime}+1}\leftarrow\mathbbm{1}_{\{s^{\prime}_{1}\}}$
@@
10:$\Delta_{N^{\prime}}(V)\leftarrow\sigma_{\hat{\mathsf{P}}^{a}_{s,N^{\prime}+1}}(V)-\frac{1}{2}\Bigl{[}\sigma_{\hat{\mathsf{P}}^{a,E}_{s,N^{\prime}+1}}(V)+\sigma_{\hat{\mathsf{P}}^{a,O}_{s,N^{\prime}+1}}(V)\Bigr{]}$
11:$\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)\leftarrow\sigma_{\hat{\mathsf{P}}^{a,1}_{s,N^{\prime}+1}}(V)+\frac{\Delta_{N^{\prime}}(V)}{\mathbb{P}(N^{\prime}=n)},\text{where }p^{\prime}(n)=\mathbb{P}(N^{\prime}=n)$ return $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$
@@
Theorem 3 (Finite Sample Complexity).
@@
The proof of Theorem 3 is in Appendix E.1, which demonstrates that setting the geometric sampling parameter to $\Psi=0.5$ ensures that the expected number of samples follows a linear growth pattern rather than an exponential one. This choice precisely cancels out the effect of the exponential sampling inherent in the truncated MLMC estimator, preventing infinite expected sample complexity. This result shows that the expected number of queries grows only linearly with $N_{\max}$, ensuring that the sampling cost remains manageable even for large truncation levels. The key factor enabling this behavior is setting the geometric distribution parameter to $0.5,$ which balances the probability mass across different truncation levels, preventing an exponential increase in sample complexity.
Theorem 4 (Exponentially Decaying Bias).
Let $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ be the estimator of ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$ obtained from Algorithm 1 the under TV uncertainty set, we have:
@@
Theorem 4 establishes that the bias of the truncated MLMC estimator decays exponentially with $N_{\max}$, ensuring that truncation does not significantly affect accuracy. This result follows from observing that the deviation introduced by truncation can be expressed as a sum of differences between support function estimates at different level, and each of which is controlled by the $\ell_{1}$-distance between transition distributions. Thus, we can use binomial concentration property to ensure the exponentially decaying bias.
The proof of Theorem 4 is in Appendix E.2. One important lemma used in the proof is the following Lemma 2, where we show the Lipschitz property for both TV and Wasserstein distance uncertainty sets.
Lemma 2.
@@
$|\sigma_{\mathcal{P}_{TV}}(V)-\sigma_{\mathcal{Q}_{TV}}(V)|\leq\|V\|_{\mathrm{sp}}\|p-q\|_{1}\;\text{and}\;|\sigma_{\mathcal{P}_{W}}(V)-\sigma_{\mathcal{Q}_{W}}(V)|\leq\|V\|_{\mathrm{sp}}\|p-q\|_{1}$ (23)
We refer the proof of Theorem 4 to Appendix E.3.
Theorem 5 (Linear Variance).
@@
Theorem 5 establishes that the variance of the truncated MLMC estimator grows linearly with $N_{\max}$, ensuring that the estimator remains stable even as the truncation level increases. The proof of Theorem 5 is in Appendix E.4, which follows from bounding the second moment of the estimator by analyzing the variance decomposition across different MLMC levels. Specifically, by expressing the estimator in terms of successive refinements of the transition model, we show that the variance accumulates additively across levels due to the binomial concentration property.
5 Robust Average Reward TD Learning
Equipped with the methods of constructing $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ for all $s\in\mathcal{S}$ and $a\in\mathcal{A}$, we now present the formal algorithm for robust policy evaluation and robust average reward for a given policy $\pi$ in Algorithm 2. Algorithm 2 presents a robust temporal difference (TD) learning method for policy evaluation in robust average-reward MDPs. This algorithm builds upon the truncated MLMC estimator (Algorithm 1) and the biased stochastic approximation framework in Section D, ensuring both efficient sample complexity and finite-time convergence guarantees.

### v3

@@ 3.2 Robust Bellman Operator
Motivated by Theorem 3.2, we define the robust Bellman operator, which forms the basis for our policy evaluation procedure.
Definition 3.3 (Robust Bellman Operator, wang2023model ).
@@
The operator $\mathbf{T}_{g}$ transforms a value function $V$ by incorporating the worst-case transition effect. A key challenge in solving the robust Bellman equation is that $\mathbf{T}_{g}$ does not satisfy contraction under standard norms, preventing the use of conventional fixed-point iteration. To cope with this problem, we establish that $\mathbf{T}_{g}$ is a contraction under some constructed semi-norm. This allows us to further develop provably efficient stochastic approximation algorithms.
@@ 4 Semi-Norm Contraction of Robust Bellman Operators
Under Assumption 3.1, we are able to establish the semi-norm contraction property. For motivation, we first establish the semi-norm contraction property of the non-robust average-reward Bellman operator for a policy $\pi$ under transition $\mathsf{P}$ defined as follows:
@@
Under ergodicity, the one‐step transition matrix (denoted as $\mathsf{P}^{\pi}$) has a unique stationary distribution $d^{\pi}$, define the stationary projector $E=\mathbf{e}^{\top}d^{\pi}$, then the fluctuation matrix (defined as $Q^{\pi}=\mathsf{P}^{\pi}-E$) has all eigenvalues strictly inside the unit circle. Standard finite‐dimensional theory (via the discrete Lyapunov equation horn2012matrix ) would produce a norm $\|\cdot\|_{Q}$ on $\mathbb{R}^{S}$ such that there is a constant $\alpha\in(0,1)$ such that for any $x\in\mathbb{R}^{S}$, $\|Q^{\pi}x\|_{Q}\leq\alpha\|x\|_{Q}$. We then build the semi-norm as follows:
@@
The concrete proof of Lemma 4.1 including the detailed construction of the semi-norm $\|\cdot\|_{\mathsf{P}}$ is in Appendix A.1, where the properties of irreducible and aperiodic finite state Markov chain are utilized. Thus, we show the (non-robust) average-reward Bellman operator $\mathbf{T}_{g}^{\mathsf{P}}$ is a strict contraction under $\|\cdot\|_{\mathsf{P}}$. Based on the above motivations, we now formally establish the contraction property of the robust average-reward Bellman operator by leveraging Lemma 4.1 and the compactness of the uncertainty sets.
Theorem 4.2.
Under Assumption 3.1, if $\mathcal{P}$ is compact, with certain restrictions on the radius of the uncertainty sets, there exists a semi-norm $\|\cdot\|_{\mathcal{P}}$ with kernel $\{c\mathbf{e}:c\in\mathbb{R}\}$ such that the robust Bellman operator $\mathbf{T}_{g}$ is a contraction. Specifically, there exist $\gamma\in(0,1)$ such that
@@
For any $\mathsf{P}\in\mathcal{P}$, the one‐step transition matrix $\mathsf{P}^{\pi}$ has a unique stationary projector $E_{\mathsf{P}}$ due to ergodicity. Since $\mathcal{P}$ is compact, the family of fluctuation matrices $\{Q_{\mathsf{P}}^{\pi}=\mathsf{P}^{\pi}-E_{\mathsf{P}}:\mathsf{P}\in\mathcal{P}\}$ has joint spectral radius strictly less than $1$. By Lemma F.1in berger1992bounded , one is able to construct an “extremal norm” (denoted as $\|\cdot\|_{\rm ext}$) under which every $Q_{\mathsf{P}}^{\pi}$ contracts by a uniform factor $\alpha\in(0,1)$. Mimicking the non-robust case in Lemma 4.1, we similarly define
@@
The concrete proof of Theorem 4.2 along with the detailed construction of the semi-norm $\|\cdot\|_{\mathcal{P}}$ and the specific radius restrictions on various uncertainty sets are in Appendix A.2. Since all the uncertainty sets listed in Section 3.1 are closed and bounded in a real vector space, these uncertainty sets are all compact and satisfy the contraction property in Theorem 4.2. We also note that the contraction factor $\gamma$ relates to the joint spectral gap of the family $\{Q^{\pi}_{\mathsf{P}}:\mathsf{P}\in\mathcal{P}\}$.
@@ 5 Efficient Estimators for Uncertainty Sets
where the learning rate $\eta_{t}$ would be specified in Section 6. The detailed analysis and complexities of the general stochastic approximation in the form of (17) is provided in Appendix B. Theorem B.1 implies that if $\hat{\mathbf{T}}_{g}(V)$, being an estimator of $\mathbf{T}_{g}(V)$, could be constructed with bounded variance and small bias, $V_{t}$ converges to a solution of the Bellman equation in (4). However, the challenge of constructing our desired $\hat{\mathbf{T}}_{g}(V)$ lies in the construction of the support function estimator $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$.
In this section, we aim to construct an estimator $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ for all $s\in\mathcal{S}$ and $a\in\mathcal{A}$ in various uncertainty sets. Recall that the support function ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$ represents the worst-case transition effect over the uncertainty set $\mathcal{P}^{a}_{s}$ as defined in the robust Bellman equation in Theorem 3.2. The explicit forms of ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$ for different uncertainty sets were characterized in (6)-(9). Our goal in this section is to construct efficient estimators $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ that approximates ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$ while maintaining controlled variance and finite sample complexity.
@@
Recall that the $\delta$-contamination uncertainty set is $\mathcal{P}^{a}_{s}=\{(1-\delta)\tilde{\mathsf{P}}^{a}_{s}+\delta q:q\in\Delta(\mathcal{S})\},$ where $0<\delta<1$ is the radius. Since the support function can be computed by (6) and the expression is linear in the nominal transition kernel $\tilde{\mathsf{P}}^{a}_{s}$. A direct approach is to use the transition to the subsequent state to construct our estimator:
@@
where $s^{\prime}$ is a subsequent state sample after $(s,a)$. Hence, the sample complexity of (18) is just one. Lemma F.3 from wang2023model states that $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ obtained by (18) is unbiased and has bounded variance as follows:
@@
Regarding TV and Wasserstein distance uncertainty sets, they have a nonlinear relationship between the nominal distribution $\tilde{\mathsf{P}}^{a}_{s}$ and the support function ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$. Previous works such as (blanchet2015unbiased, ; blanchet2019unbiased, ; wang2023model, ) have proposed a Multi-Level Monte-Carlo (MLMC) method for obtaining an unbiased estimator of ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$ with bounded variance. However, their approaches require drawing $2^{N+1}$ samples where $N$ is sampled from a geometric distribution $\mathrm{Geom}(\Psi)$ with parameter $\Psi\in(0,0.5)$. This operation would need infinite samples in expectation for obtaining each single estimator as $\mathbb{E}[2^{N+1}]=\sum^{\infty}_{N=0}2^{N+1}\Psi(1-\Psi)^{N}=\sum^{\infty}_{N=0}2\Psi(2-2\Psi)^{N}\rightarrow\infty$. To handle the above problem, we aim to provide an estimator $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ with finite sample complexity and small enough bias. We construct a truncated-MLMC estimator under geometric sampling with parameter $\Psi=0.5$ as shown in Algorithm 1.
@@
[…]
@@
10:end if
[…]
@@
Theorem 5.1 (Finite Sample Complexity).
@@
The proof of Theorem 5.1 is in Appendix C.1, which demonstrates that setting the geometric sampling parameter to $\Psi=0.5$ ensures that the expected number of samples follows a linear growth pattern rather than an exponential one. This choice precisely cancels out the effect of the exponential sampling inherent in the truncated MLMC estimator, preventing infinite expected sample complexity. This result shows that the expected number of queries grows only linearly with $N_{\max}$, ensuring that the sampling cost remains manageable even for large truncation levels. The key factor enabling this behavior is setting the geometric distribution parameter to $0.5,$ which balances the probability mass across different truncation levels, preventing an exponential increase in sample complexity.
Theorem 5.2 (Exponentially Decaying Bias).
Let $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ be the estimator of ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$ obtained from Algorithm 1 then under TV uncertainty set, we have:
@@
Theorem 5.2 establishes that the bias of the truncated MLMC estimator decays exponentially with $N_{\max}$, ensuring that truncation does not significantly affect accuracy. This result follows from observing that the deviation introduced by truncation can be expressed as a sum of differences between support function estimates at different level, and each of which is controlled by the $\ell_{1}$-distance between transition distributions. Thus, we can use binomial concentration property to ensure the exponentially decaying bias.
The proof of Theorem 5.2 is in Appendix C.2. One important lemma used in the proof is the following Lemma 5.3, where we show the Lipschitz property for both TV and Wasserstein distance uncertainty sets.
Lemma 5.3.
@@
$|\sigma_{\mathcal{P}_{TV}}(V)-\sigma_{\mathcal{Q}_{TV}}(V)|\leq(1+\frac{1}{\delta})\|V\|_{\mathrm{sp}}\|p-q\|_{1}\;\text{and}\;|\sigma_{\mathcal{P}_{W}}(V)-\sigma_{\mathcal{Q}_{W}}(V)|\leq\|V\|_{\mathrm{sp}}\|p-q\|_{1}$ (23)
We refer the proof of Theorem 5.2 to Appendix C.3.
Theorem 5.4 (Linear Variance).
@@
Theorem 5.4 establishes that the variance of the truncated MLMC estimator grows linearly with $N_{\max}$, ensuring that the estimator remains stable even as the truncation level increases. The proof of Theorem 5.4 is in Appendix C.4, which follows from bounding the second moment of the estimator by analyzing the variance decomposition across different MLMC levels. Specifically, by expressing the estimator in terms of successive refinements of the transition model, we show that the variance accumulates additively across levels due to the binomial concentration property.
6 Robust Average-Reward TD Learning
Equipped with the methods of constructing $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ for all $s\in\mathcal{S}$ and $a\in\mathcal{A}$, we now present the formal algorithm for robust policy evaluation and robust average reward for a given policy $\pi$ in Algorithm 2. Algorithm 2 presents a robust temporal difference (TD) learning method for policy evaluation in robust average-reward MDPs. This algorithm builds upon the truncated MLMC estimator (Algorithm 1) and the biased stochastic approximation framework in Section B, ensuring both efficient sample complexity and finite-time convergence guarantees.